OpenStudy (anonymous):

Solve the equation. (Enter your answers as a comma-separated list.) square root of 12x +24 = 6, square root of x

OpenStudy (anonymous):

\[\sqrt{12x+24}= 6 \sqrt{x}\]

OpenStudy (jamesj):

Step 1 always with square roots is to get rid of them. So square both sides and see what you get

OpenStudy (anonymous):

the answer I got was 12/11 but it was wrong

OpenStudy (jamesj):

What is the equation you get upon squaring both sides?

OpenStudy (anonymous):

12x +24 = 36 +x

OpenStudy (jamesj):

No. 12x + 24 = 36x as \( (6\sqrt{x})^2 = 36 x \)

OpenStudy (jamesj):

\[ (6\sqrt{x})^2 = 6^2 \sqrt{x}^2 = 36 x \]

OpenStudy (anonymous):

thank you so much